The length of a rectangular lot is 4m (meters) longer than the width. If the area of the lot is 12m^2, what are its dimensions?
Answer by Cromlix(4381)

Hi there,
Make width = 'x'
Length = 4 + x
Area of a rectangle
= Length x Width
12m^2 = (4 + x)x
12m^2 = 4x + x^2
Put in ax^2 + bx^2 + c = 0 form
x^2 + 4x - 12 = 0
Factorise
(x + 6)(x - 2)= 0
x + 6 = 0
x= -6 (No answer as -ve)
x - 2 = 0
x = 2
Therefore:
Width = 2m
Length = 6m
Hope this helps:-)